### v1

@@ 3.2.2 Matching annotations and OCR
    Questions are generated using the template What is the [key name]? (e.g., What is the Address?). For XFUND, this template was translated to match document languages. Datasets with pre-defined questions (DUDE, MP-DocVQA, SP-DocVQA) used their own questions.
@@ 3.2.3 Rephrasing questions
After completing the matching between annotations and OCR, the questions for the new dataset are generated. Inspection of these questions, which followed a simple template-based structure, revealed that they are often grammatically incorrect, overly simplistic, and consistently adhered to the same pattern. This raised concerns that training an LLM on these questions could introduce bias, potentially leading to poor performance on questions written by humans, which may not follow the template.
To mitigate this issue, we employed the Mistral 7B model [36] to correct and rewrite the questions, aiming to fix errors and introduce linguistic diversity. Other Mistral models, such as Mistral Large [37] and Mixtral 8x7B [38], were also tested, but they produced overly complex, verbose, and unnatural questions.
@@
This procedure was applied to most questions in the dataset, adding a new attribute, rephrased_question. Questions from DUDE, MP-DocVQA, and SP-DocVQA were excluded as they were already human-written. Additionally, questions from XFUND were excluded due to concerns over the model’s ability to generate questions in languages other than English.
@@
Table 2: Overall dataset statistics.
In Figure 2 you can observe an example of the final format of the dataset questions, including the rephrased version of the questions.
@@ 3.3 Statistics & splits
The dataset is split into training, validation, and test sets, using an 80-10-10 split based on document count, where all questions related to a single document are contained within the same set. Table 2 gives an overview of the dataset’s size and sources distribution. Detailed statistics can be found in the Supplementary Material.
To ensure that question types and document layouts are uniformly distributed across the three sets, documents from each source dataset are sampled separately. Specifically, documents from Deepform are split in an 80-10-10 ratio, followed by documents from FATURA, DUDE, and all the others. The union of these individual splits yields the final training, validation, and test sets, with balanced layout and document types across all sets.
Some of the pages annotated using the proposed algorithm and belonging to BoundingDocs are shown in figures 3 and 4. For illustration purpose colored rectangle is drawn around the fields corresponding to the correct answers to the questions.
@@ 3.4 Dataset examples
Figure 5: Finetuning and testing pipeline. The rectangles represent processes while the parallelograms represent foundation LLM models.
In Fig. 3 (Deepform) and Fig. 4 (VRDU Registration Form) it is possible to observe two pages while Table 3 contains their QA pairs. In Fig. 3 the extracted fields are the advertiser’s name and the gross amount for the various transmissions. In Fig. 4 the fields to be extracted are only two: the registrant name and the registration number.
These two examples illustrate how, despite the high number of documents in the collection, the potential amount of information present in the documents is underutilized, as the annotated fields are few compared to the entire body of the documents, indicating that the potential of this large document collection is not being properly exploited.
@@ 4 Experimental Results
Table 4 presents our experimental results across the different datasets and model configurations. The finetuning and testing pipeline implemented is summarized and plotted in Figure 5.
Model Deepform DUDE FATURA FUNSD XFUND SP-VQA
Mistral-7B-v0.3$\star$ 42.3 9.1 6.8 14.3 6.1 22.2
0.22% 16.19% 1.03% 1.23% 15.85% 10.00%
[…]
0.47% 5.52% 0.12% 3.68% 9.55% 2.50%
[…]
6.79% 64.76% 7.40% 5.52% 2.07% 52.50%
Template-Template 97.7 70.5 99.9 75.7 70.1 75.3
0.00% 0.00% 0.00% 0.61% 0.61% 0.00%
Template-Rephrased 96.8 70.9 91.5 71.1 68.8 70.2
3.75% 1.14% 0.23% 0.00% 1.53% 2.50%
Rephrased-Template 97.7 70.4 99.7 71.8 67.6 76.6
0.00% 0.00% 0.00% 0.00% 0.67% 0.00%
Rephrased-Rephrased 97.1 71.2 99.8 72.3 68.2 76.1
0.00% 0.00% 0.00% 0.00% 0.92% 0.00%
Reph.-Reph.-bbox 97.7 73.4 99.3 78.8 71.2 82.1
5.01% 4.95% 5.97% 17.79% 10.34% 5.00%
Reph.-Reph.-bbox 97.7 72.1 99.3 74.7 70.3 83.0
w/ regex 0.80% 0.38% 4.50% 4.29% 0.98% 0.00%
Model Kl. Charity Kl. NDA MP-VQA VRDU-Ad VRDU-Reg. W. Avg.
Mistral-7B-v0.3$\star$ 21.2 32.5 12.5 23.1 28.6 22.4
4.84% 10.45% 7.86% 0.22% 1.52% 3.32%
[…]
2.93% 6.72% 3.54% 0.65% 6.09% 1.77%
[…]
2.07% 0.00% 52.82% 26.80% 4.31% 20.37%
Template-Template 91.9 66.3 75.5 96.7 96.5 91.3
0.00% 0.00% 0.06% 0.00% 0.00% 0.04%
Template-Rephrased 92.5 63.7 73.6 87.1 96.0 87.8
0.40% 1.49% 1.03% 0.13% 0.00% 1.62%
Rephrased-Template 91.8 61.1 73.6 96.2 95.2 90.7
0.00% 0.00% 0.06% 0.00% 0.00% 0.04%
Rephrased-Rephrased 92.3 64.4 73.3 96.4 95.7 90.6
0.00% 0.00% 0.07% 0.00% 0.00% 0.04%
Reph.-Reph.-bbox 92.8 61.6 76.0 96.4 96.7 91.6
4.34% 0.75% 7.16% 1.47% 0.51% 5.64%
Reph.-Reph.-bbox 92.9 61.9 75.8 96.1 96.8 91.3
w/ regex 0.03% 0.00% 0.35% 0.26% 0.00% 1.53%
Table 4: ANLS* scores and JSON parsing error percentages across datasets for each model in our custom dataset. ANLS* scores measure accuracy in answering document questions, while the bottom value in each cell shows JSON parsing errors, indicating output consistency. The first three rows list instruct models ($\star$); all remaining rows are fine-tuned versions of Mistral-7B-v0.3. Model names follow the ‘[training question type]-[testing question type]‘ format (e.g., ‘Template-Rephrased‘ means trained on template questions, tested on rephrased ones). ”bbox” indicates layout information is included in the prompt, and ”w/ regex” denotes that values were extracted with regex if JSON parsing failed. The ”W. Avg” column provides a weighted average across datasets, with bold and underlined values marking the top two scores per dataset.
@@ 4.1 Evaluation Metrics
We use the standard metric ANLS* [39], which supports a wider range of tasks including line-item extraction and document-processing tasks.
For each model-dataset pair in our results, we report two key measurements: the ANLS* value (rescaled between 0 and 100 for easier reading) and the percentage of non-JSON parsable responses relative to the total number of queries. The weighted average provides a comprehensive overview based on the number of examples for each dataset. For ANLS*, higher values indicate better performance, while for non-parsable responses, lower percentages are preferable. In our results presentation, the best values for each dataset are bolded, and second-best values are underlined.
@@ 4.2 Prompt Construction
In this study, each question in the dataset may have answers located on multiple pages. The significant computational costs associated with multi-page document processing, as reported e.g. by Multi PageDocVQA [15], together with the context size limitation of smaller LLMs, make us opt for an atomic approach instead of encoding everything in a single prompt.
For questions requiring information from multiple pages, we generate independent prompts for each relevant page, appending the same question to each prompt. For instance, if a five-page document contains relevant information on pages 2 and 4, we generate two prompts—one containing page 2’s content and the other page 4’s—each coupled with the question.
Each prompt includes the document text, the question, and a specification for the answer format (JSON), facilitating structured data extraction.
@@ 4.3 Baseline Models
We evaluated three popular open-weight models as baselines: Mistral 7B Instruct v0.3 [36], Llama 3 8B Instruct [40], and Phi 3.5 3.8B Instruct [41]. These models were chosen for their established performance and recognition in the NLP community. Testing was conducted on our custom dataset to establish initial benchmarks.
@@ 4.4 Ablation Study: Question Formulation
For investigating the impact of question formulation, we selected the Mistral 7B v0.3 (base version) for fine-tuning. We evaluated two types of questions—template-based (simple, consistent format) and rephrased (more varied, user-friendly language). Each model was tested with both question types, resulting in four experimental conditions:

### v2

@@ 3.2.3 Questions formulation
    Questions are generated using the template What is the [key name]? (e.g., What is the Address?). Note that [key name] does not refer to the actual name assigned to the key by the datasets (e.g., program_desc) but rather to a refined, natural-language version defined by us (e.g., program description). We manually refined around 10 keys per dataset. This ensures that questions are always in natural language—though not necessarily correct—but consistently composed of meaningful words. For XFUND, the question template was automatically translated to match document languages. Datasets with pre-defined questions (DUDE, MP-DocVQA, SP-DocVQA) used their own questions.
@@ 3.2.4 Rephrasing questions
After the previously described steps, the questions for the new dataset are generated. Inspection of these questions, which followed a simple template-based structure, revealed that they are often grammatically incorrect, overly simplistic, and consistently adhered to the same pattern. This raised concerns that fine-tuning an LLM on these questions could introduce bias, potentially leading to poor performance on questions written by humans, which may not follow the template.
To mitigate this issue, we employed the Mistral 7B model [jiang2023mistral] to correct and rewrite the questions, aiming to fix errors and introduce linguistic diversity. Other Mistral models, such as Mistral Large [mistral2024large] and Mixtral 8x7B [mistral2023mixtral], were also tested, but they produced overly complex, verbose, and unnatural questions.
@@
This procedure was applied to most questions in the dataset, adding a new attribute, rephrased_question. Questions from DUDE, MP-DocVQA, and SP-DocVQA were excluded as they were already human-written.
To ensure quality and limit hallucinations, we iteratively refined the prompt during the design phase and validated the outputs through manual sampling. Additionally, the LLM was instructed to preserve the semantic meaning a from the template questions, using the original answer as a starting point.
@@
Table 2: Overall statistics of BoundingDocs, divided by source.
In Figure 2 we show one example of the final format of the dataset questions, including the rephrased version of the questions.
@@ 3.3 Statistics & splits
The dataset is split into training, validation, and test sets using an 80-10-10 split based on document count, ensuring that all questions related to a single document remain within the same set. Since the dataset comprises diverse sources with varying structures and question types, we propose our own split to ensure a balanced distribution of document layouts and question types across all sets. Table 2 provides an overview of the dataset’s size and source distribution, while detailed statistics can be found in the Supplementary Material.
To achieve this balance, documents from each source dataset are sampled separately. Specifically, documents from Deepform are split in an 80-10-10 ratio, followed by documents from FATURA, DUDE, and all other sources. The union of these individual splits yields the final training, validation, and test sets, ensuring that all sets reflect the dataset’s overall diversity.
Some of the pages annotated using the proposed algorithm and belonging to BoundingDocs are shown in Figures 3 and 4. For illustration purpose, colored rectangles are drawn around the fields corresponding to the correct answers to the questions.
@@
It should be noted, as previously mentioned, that for datasets such as Kleister, DUDE, and DocVQA, the test splits are not publicly available. Consequently, documents from these test splits are not included in our dataset. For datasets like FATURA and Deepform, which do not have predefined splits, this issue does not arise. However, further discussion is needed regarding FUNSD, XFUND, and VRDU, as they do have defined splits. In the case of FUNSD and XFUND, a portion of the documents from their test splits were included in the random process used to define the splits in BoundingDocs. As a result, some of these documents are part of the BoundingDocs training split. Specifically, out of the $38,515$ documents in the training split, $313$ originate from the FUNSD/XFUND test sets, accounting for approximately $0.8\%$. Regarding VRDU, its repository provides multiple splits depending on the task of interest, meaning there is no single definitive test split. Therefore, VRDU has been treated similarly to FATURA and Deepform. This analysis informs users that incorporating BoundingDocs into training data does not prevent models from being evaluated on benchmarks like DUDE, DocVQA, or Kleister. This leads to the conclusion that BoundingDocs can serve a dual purpose: it can be used as a pretraining tool without significant risk of contamination with established benchmarks, allowing these benchmarks to be employed for evaluating the resulting models. Additionally, BoundingDocs can function as a benchmark itself for assessing models in a question answering setting, even for datasets and document types that were not originally designed for this task.
3.4 Annotation assessment
To provide insights into the dataset quality, we analyzed the frequency of multiple matches within a single page, as mentioned in Section 3.2.2. We found that in approximately 20% of cases, the annotated value appeared more than once on the same page. However, this does not necessarily indicate that 20% of the dataset contains false positives. To further investigate this phenomenon, we conducted a manual audit of annotations with multiple matches. We randomly sampled 20 pages containing at least one question with multiple valid answers, selecting one such question per page. We successfully sampled 20 pages for most datasets, with the exception of FUNSD (19 pages), SP-DocVQA (13 pages), and FATURA and VRDU Registration Form (0 pages for both). For each sampled page-question pair, we manually examined all annotated answer bounding boxes to verify whether all instances referred to the same logical answer or if some were unrelated occurrences of the same textual value. An annotation was considered correct only if all annotated answers corresponded to valid responses to the question, rather than being semantically unrelated occurrences of the same text. The results, reported in Table 3, reveal that approximately 38% of questions with multiple answers per page contain at least one answer that is not logically related to the question. Extrapolating from this finding, we estimate that approximately 7% of all annotations (38% of the 20% with multiple matches) may exhibit this issue. Examining the results by data type, we observe distinct patterns in annotation accuracy. Acronyms achieve perfect accuracy (100%), as they are highly specific and rarely repeated for different purposes within documents. Named entities and currency values show moderate accuracy (68.12% and 70.00%, respectively), while dates perform slightly lower at 66.67%. Numbers exhibit the lowest accuracy rate (46.88%), as numerical values frequently appear in unrelated sections of documents such as page numbers, reference codes, or unrelated quantities. The “Other” category, which includes miscellaneous data types, shows an accuracy of 40.00%, indicating challenges in handling diverse or less structured information.
Data Type Total %
Currency 20 14 70.00
Date 18 12 66.67
Number 32 15 46.88
Acronym 8 8 100.00
Named entities 69 47 68.12
Other 25 10 40.00
Total 172 106 61.63
Table 3: Annotation quality statistics by data type for questions with multiple valid answers per page.
@@ 3.5 Dataset examples
Figure 5: Experimental framework showing the different prompting strategies implemented for vision and text language models. Vision LLMs (blue) are prompted with images and rephrased questions, while text LLMs (red) are prompted with three different configurations of page content and questions. Some models underwent supervised fine-tuning before testing, while others were evaluated directly.
In Fig. 3 (Deepform) and Fig. 4 (VRDU Registration Form) it is possible to analyze two pages the corresponding QA pairs are shown in Table 4. In Fig. 3 the extracted fields are the advertiser’s name and the gross amount for the various transmissions. In Fig. 4 the fields to be extracted are only two: the registrant name and the registration number.
These two examples illustrate how, despite the large number of documents in the collection, the potential amount of information present in the documents is underutilized, as the annotated fields are few compared to the entire body of the documents, indicating that the potential of this large document collection is not being properly exploited.
To provide a concrete response to the issue raised in Section 3.2.2, as shown in Fig. 3, the value \$119,000.00 has been detected and annotated four times on the page. The two entries on the left part of the page are certainly correct, as they refer to totals: the first is related to the date and the second to the TV transmission. Regarding the values stacked on the right side of the page, the value beneath represents the sum of all those above it, and in this case, the numbers match. This validates the annotation, as the total refers to the same entity. If there were additional entries, only the total would be annotated, as the partial values would necessarily differ.
@@ 4 Experimental results
Tables 5, 6, and 7 present our experimental results across the different datasets and model configurations. The fine-tuning and testing pipeline implemented is summarized and plotted in Figure 5, which depicts the various types of prompts used depending on the model type: for vision models, we provided the image along with the question only, while for traditional LLMs, we used the OCR-extracted page text, the question, and, for some ablation, the bounding boxes of the text. Notably, all the fine-tuning experiments described were conducted using only 10% of the training split and evaluated on the proposed test split.
The three tables follow the same format, where the columns correspond to the documents in the BoundingDocs test set derived from the respective dataset. All models and configurations have been evaluated on the proposed test split. It is crucial to emphasize that these results are not directly comparable to established benchmarks (e.g., DocVQA or DUDE) or to performances reported in other studies using the same datasets. Additionally, comparisons between different model types (e.g., LLMs vs. vision LLMs) should be avoided due to differences in prompting techniques and LoRA fine-tuning parameters (details in the Supplementary Material). The test split is specifically determined by our question-filtering strategy and dataset construction methodology. Therefore, these results primarily serve to illustrate the dataset’s difficulty and provide a baseline, rather than to position our approach relative to the state of the art.
@@ 4.1 Evaluation metrics
We use the standard metric ANLS* [peer2024anlsuniversaldocument], which supports a wider range of tasks including line-item extraction and document-processing tasks.
For each model-dataset pair in our results, we report two key measurements: the ANLS* value (rescaled between 0 and 100 for easier reading) and the percentage of non-JSON parsable responses relative to the total number of queries. The weighted average provides a comprehensive overview based on the number of examples for each dataset. For ANLS*, higher values indicate better performance, while for non-parsable responses, lower percentages are preferable. In the tables, the best values for each dataset are in bold, whike the second-best values are underlined.
@@ 4.2 Prompt construction
In this study, each question in the dataset may have answers distributed across multiple pages. Processing multi-page documents poses significant computational challenges, as noted in works such as Multi-PageDocVQA [TITO2023109834]. Additionally, the context size limitations of smaller LLMs make encoding all pages into a single prompt impractical. To address these constraints, we adopt an atomic approach, processing each relevant page separately rather than constructing a single comprehensive prompt.
Our experiments follow the ‘oracle’ setup described in [TITO2023109834], where only the page containing the answer is provided as input to the model. This setup isolates the model’s answering capabilities from variations in input sequence length, serving as a theoretical upper bound on performance—assuming the method correctly identifies the relevant page. As a result, these baselines reflect the intrinsic difficulty of the questions and document layouts rather than external factors such as content filtering. Nevertheless, the dataset includes complete documents, allowing users to implement and evaluate multi-page methods, which must also handle the task of identifying relevant pages. For questions requiring information from multiple pages, we generate independent prompts for each relevant page, appending the same question to each. For instance, if a five-page document contains relevant information on pages 2 and 4, we create two separate prompts: one incorporating content in page 2 and the other incorporating page 4, both paired with the same question.
Each prompt consists of three components: the document text, the question, and a given answer format (JSON) to facilitate structured data extraction.
@@ 4.3 Baseline models
Model
Deepform
DUDE
FATURA
FUNSD
XFUND
SP-VQA
Kl. Char.
Kl. NDA
MP-VQA
VRDU Ad
VRDU Reg
W. Avg.
Mistral-7B-v0.3 42.3 9.1 6.8 14.3 6.1 22.2 21.2 32.5 12.5 23.1 28.6 22.4
0.22% 16.19% 1.03% 1.23% 15.85% 10.00% 4.84% 10.45% 7.86% 0.22% 1.52% 3.32%
Llama-3-8B 83.9 60.0 35.6 70.5 38.4 73.7 72.8 25.3 62.4 71.2 37.9 62.9
0.47% 5.52% 0.12% 3.68% 9.55% 2.50% 2.93% 6.72% 3.54% 0.65% 6.09% 1.77%
Phi-3.5-3.8B 66.4 45.2 24.7 55.8 51.3 50.2 63.9 48.1 54.4 59.2 57.9 51.6
6.79% 64.76% 7.40% 5.52% 2.07% 52.50% 2.07% 0.00% 52.82% 26.80% 4.31% 20.37%
Claude 3.7 Sonnet 62.2 72.4 52.5 78.6 54.2 85.8 76.8 16.6 76.9 90.6 79.6 65.7
0.00% 0.00% 0.00% 1.22% 0.00% 0.00% 0.00% 0.00% 0.04% 0.00% 0.00% 0.02%
Qwen2-VL-2B 35.3 60.0 36.1 71.6 48.7 88.3 48.7 28.5 68.1 52.3 62.3 45.8
0.57% 0.38% 0.05% 1.23% 1.41% 0.00% 0.29% 1.49% 1.21% 1.25% 0.00% 0.54%
Qwen2-VL-7B 44.7 69.3 47.5 73.2 59.7 84.5 70.4 20.1 72.9 56.1 70.0 55.6
0.04% 0.00% 3.44% 0.00% 0.37% 0.00% 0.11% 0.00% 0.14% 0.09% 0.00% 0.93%
Table 5: Baseline models. ANLS* scores and JSON parsing error percentages across datasets for baseline models on our test split. ANLS* scores measure accuracy in answering document questions, while the bottom value in each cell shows JSON parsing errors, indicating output consistency. The ”W. Avg” column provides a weighted average across datasets, with bold and underlined values marking the top two scores per dataset.
We evaluated three popular open-weight models as baselines: Mistral 7B Instruct v0.3 [jiang2023mistral], Llama 3 8B Instruct [touvron2023llama], and Phi 3.5 3.8B Instruct [abdin2024phi3]. These models were chosen for their established performance and recognition in the NLP community.
In addition to these text-only models, we also tested three multimodal models capable of processing both text and images: Claude 3.7 Sonnet [claudesonnet] and Qwen2-VL [DBLP:journals/corr/abs-2409-12191] in its 2B and 7B variants. This selection ensures a diverse range of model types and sizes, allowing for a comprehensive evaluation. These tests served to establish an initial benchmark and are reported in Table 5.
Through a qualitative analysis based on the results of our preliminary experiments and observations of dataset examples, we identified the most complex layouts and the sources for which the proposed task is particularly challenging. Notably, Deepform and VRDU Ad Buy Form contain numerous similar values that can correspond to different entries in the financial report, complicating information extraction. Despite its seemingly simple layout, Kleister NDA also presents significant challenges. This suggests that the lack of a clear structure — characteristic of contract pages with dense and compact text — impedes the models’ ability to retrieve highly specific information. A comprehensive overview of the various sources that compose BoundingDocs can be found in the Supplementary Material.
@@ 4.4 Ablation study: question formulation
Model
Deepform
DUDE
FATURA
FUNSD
XFUND
SP-VQA
Kl. Char.
Kl. NDA
MP-VQA
VRDU Ad
VRDU Reg
W. Avg.
Templ.-Templ. 97.7 70.5 99.9 75.7 70.1 75.3 91.9 66.3 75.5 96.7 96.5 91.3
0.00% 0.00% 0.00% 0.61% 0.61% 0.00% 0.00% 0.00% 0.06% 0.00% 0.00% 0.04%
Templ.-Reph. 96.8 70.9 91.5 71.1 68.8 70.2 92.5 63.7 73.6 87.1 96.0 87.8
3.75% 1.14% 0.23% 0.00% 1.53% 2.50% 0.40% 1.49% 1.03% 0.13% 0.00% 1.62%
Reph.-Templ. 97.7 70.4 99.7 71.8 67.6 76.6 91.8 61.1 73.6 96.2 95.2 90.7
0.00% 0.00% 0.00% 0.00% 0.67% 0.00% 0.00% 0.00% 0.06% 0.00% 0.00% 0.04%
Reph.-Reph. 97.1 71.2 99.8 72.3 68.2 76.1 92.3 64.4 73.3 96.4 95.7 90.6
0.00% 0.00% 0.00% 0.00% 0.92% 0.00% 0.00% 0.00% 0.07% 0.00% 0.00% 0.04%
Reph.-Reph.-bbox 97.7 73.4 99.3 78.8 71.2 82.1 92.8 61.6 76.0 96.4 96.7 91.6
5.01% 4.95% 5.97% 17.79% 10.34% 5.00% 4.34% 0.75% 7.16% 1.47% 0.51% 5.64%
Reph.-Reph.-bbox 97.7 72.1 99.3 74.7 70.3 83.0 92.9 61.9 75.8 96.1 96.8 91.3
w/ regex 0.80% 0.38% 4.50% 4.29% 0.98% 0.00% 0.03% 0.00% 0.35% 0.26% 0.00% 1.53%
Table 6: Ablation study. ANLS* scores and JSON parsing error percentages across datasets for each prompting configuration on our test split. ANLS* scores measure accuracy in answering document questions, while the bottom value in each cell shows JSON parsing errors, indicating output consistency. The ”W. Avg” column provides a weighted average across datasets, with bold and underlined values marking the top two scores per dataset.
Model
Deepform
DUDE
FATURA
FUNSD
XFUND
SP-VQA
Kl. NDA
MP-VQA
VRDU Ad
W. Avg.
Templ.-Templ. 97.7 70.5 99.9 75.7 70.1 75.3 66.3 75.5 96.7 91.3
0.00% 0.00% 0.00% 0.61% 0.61% 0.00% 0.00% 0.06% 0.00% 0.04%
Templ.-Reph. 96.8 70.9 91.5 71.1 68.8 70.2 63.7 73.6 87.1 87.8
3.75% 1.14% 0.23% 0.00% 1.53% 2.50% 1.49% 1.03% 0.13% 1.62%
Reph.-Templ. 97.7 70.4 99.7 71.8 67.6 76.6 61.1 73.6 96.2 90.7
0.00% 0.00% 0.00% 0.00% 0.67% 0.00% 0.00% 0.06% 0.00% 0.04%
Reph.-Reph. 97.1 71.2 99.8 72.3 68.2 76.1 64.4 73.3 96.4 90.6
0.00% 0.00% 0.00% 0.00% 0.92% 0.00% 0.00% 0.07% 0.00% 0.04%
Reph.-Reph. 97.7 73.4 99.3 78.8 71.2 82.1 61.6 76.0 96.4 91.6
+bbox 5.01% 4.95% 5.97% 17.79% 10.34% 5.00% 0.75% 7.16% 1.47% 5.64%
For investigating the impact of question formulation, we selected the Mistral 7B v0.3 (base version) for fine-tuning (details in the Supplementary Material). Results are reported in Table 6. We evaluated two types of questions—template-based (simple, consistent format) and rephrased (more varied, user-friendly language). Each model was tested with both question types, resulting in four experimental conditions: